# Comparison Performance of Spectrogram and Scalogram as Input of Acoustic Recognition Task

### v3

@@ Abstract
Index Terms:
spectrogram and scalogram, acoustic recognition, Convolutional Neural Networks
@@ I Introduction
In the adoption of spectrograms as images for speech emotion recognition, Convolutional Neural Networks (CNNs) were employed to predict emotions in audio speech, as demonstrated in the study by [1]. The model, comprising only three convolutional layers and three fully connected layers, outperformed the well-known pre-trained model AlexNet. A similar approach using CNNs and spectrograms was proposed in the study by [2] for acoustic event detection (AED) on audio recordings from a surveillance system. This method outperformed a benchmark AED-capable system. The study by [3] investigated abnormal machinery sound detection using spectrograms and an autoencoder. Two types of audio – normal and abnormal sounds collected in a factory – were transformed into spectrograms, which were then used for classification tasks to determine whether the machinery was in normal condition or damaged.
@@
Scalograms, another audio feature extraction method, have seen increasing application in recent years. An implementation of CNNs to detect milling chatter using scalograms was developed in the study by [4]. Their system, designed for real-time detection, achieved superior performance compared to benchmark methods utilizing traditional machine learning. A similar approach to domestic audio classification was applied in the study by [5]. Scalograms were produced using the output of the continuous Wavelet Transform (CWT) and then fed into a pre-trained neural network to predict whether the sound originated from social activities or a vacuum cleaner. Utilizing scalograms and CNNs for audio scene modeling, the research by [6] demonstrated better performance for the model trained with scalograms compared to one using spectrograms.
@@
As illustrated above, while numerous studies have explored audio recognition using spectrograms or scalograms in conjunction with deep learning, few have directly compared the performance of these two feature extraction methods. Existing benchmarks often offer indirect comparisons, as they involve different prediction models rather than a comprehensive evaluation of the feature extractors themselves, as seen in the study by [6]. Or the study by [7] compared different transforms but focused solely on predictive performance, lacking a comparative analysis of how the attributes of the transforms influence predictive performance and computational expense. The research by [8] compared various time-frequency transforms but used discrete coefficients directly instead of scalograms or spectrograms, limiting the models’ performance. Given these limitations, this paper aims to develop an approach for a rigorous comparison of spectrograms and scalograms as audio feature extractors. To this end, the experimental design for both methods is kept identical, including the original audio data used, the format of input data, the configuration of the CNNs models, and the training and evaluation methods. The sole distinction in this study is the type of feature extractor: spectrogram versus scalogram.
@@ II-A Short-time Fourier transform
The STFT [9] is a method used to analyze signals by dividing them into segments based on time frames and performing Fourier transforms (FT) on each segment, denoted as (1). This approach allows the signal to be examined in both the time shift $\tau$ and frequency shift $\omega$ simultaneously. By applying a windowing function, denoted as $\gamma(t)$, to isolate specific portions of the signal, STFT computes the spectrum over short durations, enabling it to capture not only the frequencies present in the signal but also how these frequencies change over time.
[…]
@@
In practical applications, signals are processed in discrete time formats, and the STFT is computed using a window of length N, referred to as the frame size, denoted as (2). Since shifting sample-wise results in redundant data due to the similarity of neighboring spectra, a larger step size, known as the hop size H, is used, typically set to half the window length H=N/2. This choice strikes a balance between achieving high time resolution and managing data volume effectively [9]. Each sample point in the data sequence, where STFT is calculated, corresponds to a frame index m defined as m=M/H, with M representing the total data length. Within each frame, computing the FT of a discrete signal of N points in time results in a frequency vector of N points. Due to the symmetric nature of the Fourier spectrum, only the first half of the spectrum (k=0 to N/2) is typically considered to avoid redundancy. Consequently, the output of STFT is structured as a matrix with dimensions (M/H, N/2+1) where each element X(m,k) represents the complex Fourier coefficient of the $m^{th}$ time frame at frequency index k.
To enhance the visualization of STFT results, spectrograms are commonly employed [10]. Represented as a type of heatmap image, they provide a two-dimensional representation of signal energy across time and frequency dimensions. Time is arranged along the x-axis, and frequency along the y-axis. The intensity of color in each pixel of the spectrogram corresponds to the degree of signal energy. For instance, the colors range from dark blue, indicating lower energy, to dark red, indicating higher energy, as illustrated in Fig. 1(a).
@@ II-B Wavelet transform
WT is a technique used to transform a signal into a representation that enhances specific features for further processing [11]. It decomposes a signal from the time domain into the time-frequency domain using a mother wavelet, denoted as (3). The WT depends on parameters of time shift b and frequency scale a. The normalization factor 1/$\sqrt{\textit{a}}$ ensures energy normalization across scales. During transformation, the mother wavelet $\psi(t)$ undergoes dilation and contraction as the scale a varies, and then slides along the time axis with step b to convolve with the signal x(t), generating a matrix of coefficients.
@@ II-C Time and frequency resolution of transforms
To achieve a comprehensive analysis of a signal, both time resolution and frequency resolution should be maximized. This implies capturing signal variations with minimal shifts in both time and frequency. However, the uncertainty principle [10] imposes a constraint on the tradeoff between time and frequency resolution, as described in (5). Utilizing a narrow time window (small $\Delta_{t}$) allows for high time resolution in STFT or WT, but results in poor frequency resolution (large $\Delta_{\omega}$). Conversely, employing a large time window yields poor time resolution but improved frequency resolution.
@@ III-A Workflow
The entire workflow of the experiment in this study is presented in Fig. 3. Initially, the audio data is normalized to constrain the amplitude values within the range (0, 1) before being transformed using the STFT and WT, respectively. Following the transformations, the matrices produced as outputs from the STFT and WT are used to create spectrograms and scalograms. These are then plotted as images and split into training and validation sets for a CNNs model. The performance of the trained CNNs is then recorded and compared between the two types of data, using metric the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [12].
@@ III-B MIMII audio dataset
The acoustic recognition task in this study is conducted using the MIMII audio dataset for machinery fault detection [3]. This dataset was created to provide real-life sound data from factories for various machine types, including fans, pumps, sliders, and valves. These machines produce both stationary and non-stationary sounds, presenting varying levels of difficulty in distinguishing between normal and abnormal conditions. Background noise from various factory environments is recorded and mixed with the target machine sounds to simulate real-world conditions at three levels of signal-to-noise ratio (SNR): -6 dB, 0 dB, and 6 dB. Each audio recording has a length of 10 seconds, sampled at 16 kHz, resulting in each discrete audio 160,000 samples.
@@ III-D Implementation of Short-Time Fourier Transform
The STFT is conducted in a time-discrete manner using Librosa [13]. Initially, the frame size N and hop size H are treated as hyperparameters aimed at optimizing the performance of the trained model during the initial stages of training. Subsequently, a configuration of STFT with N = 1024 and H = 512 demonstrating good predictive performance is selected, resulting a matrix with dimensions (513, 313). This configuration is then applied uniformly across the entire audio dataset to generate the STFT representations.
@@ III-E Implementation of Continuous Wavelet Transform
The CWT is also executed in a time-discrete manner using Pywavelets [14]. The translation is performed sample-wise, and the scale parameter encompasses a range of consecutive natural numbers. Given that the scale determines the frequency resolution, various scales are evaluated during the training of the CNNs model. Then the scale from 2 to 129, demonstrating optimal predictive performance, is selected, producing output matrices with dimensions (16000, 128). Subsequently, this selected scale is employed to generate scalograms for the entire audio dataset.
@@ III-F Implementation of Convolutional Neural Networks
The primary focus of this study is to compare the performance between spectrograms and scalograms. Therefore, the CNNs model is designed in an exploratory manner, ensuring it achieves satisfactory performance. The Tensorflow [15] is used for CNNs implementation. Several configurations of CNNs are evaluated, and an effective configuration is chosen based on its performance, as outlined in Table I.
@@ IV-A Benchmark study
The benchmark study [16], representing the state-of-the-art for supervised methods on the same dataset, employed Mel-frequency cepstral coefficients (MFCCs) as spectral features. Unlike the approach of transforming data into compressed images, this model directly utilized tuples generated by MFCCs. Prior to being used for training, the tuples were normalized to have zero mean and unit standard deviation. The prediction model used in this study was a Multilayer Perceptron (MLP). Additionally, the tuples of abnormal sounds were oversampled to achieve a balanced dataset. Detailed comparisons are presented in Table II.
@@ IV-B Performance evaluation
For the comparison between spectrogram and scalogram within each machine type, the results are presented in Table V and visualized in Figure 6. Overall, spectrograms consistently outperform scalograms, with the exception of one case involving valves where CNNs predicts audio based on scalograms more effectively than those based on spectrograms. This exception arises due to the non-stationary nature of valve audio signals [3], which are impulsive and sparse in time. The WT used in scalogram generation leverages its multiresolution feature to enhance feature extraction for such non-stationary signals [10]. Conversely, when evaluating performance on fan audio signals, spectrograms achieve significantly higher performance compared to scalograms. This difference can be attributed to the stationary nature of fan sounds [3], where the constant time and frequency resolution of spectrograms are more adept at feature extraction than scalograms.

### v4

@@ I Introduction
In the adoption of spectrograms as images for speech emotion recognition, Convolutional Neural Networks (CNNs) were employed to predict emotions in audio speech, as demonstrated in the study by [badshah2017speech]. The model, comprising only three convolutional layers and three fully connected layers, outperformed the well-known pre-trained model AlexNet. A similar approach using CNNs and spectrograms was proposed in the study by [santos2021audio] for acoustic event detection (AED) on audio recordings from a surveillance system. This method outperformed a benchmark AED-capable system. The study by [purohit2019mimii] investigated abnormal machinery sound detection using spectrograms and an autoencoder. Two types of audio – normal and abnormal sounds collected in a factory – were transformed into spectrograms, which were then used for classification tasks to determine whether the machinery was in normal condition or damaged.
@@
Scalograms, another audio feature extraction method, have seen increasing application in recent years. An implementation of CNNs to detect milling chatter using scalograms was developed in the study by [tran2020milling]. Their system, designed for real-time detection, achieved superior performance compared to benchmark methods utilizing traditional machine learning. A similar approach to domestic audio classification was applied in the study by [copiaco2019scalogram]. Scalograms were produced using the output of the continuous Wavelet Transform (CWT) and then fed into a pre-trained neural network to predict whether the sound originated from social activities or a vacuum cleaner. Utilizing scalograms and CNNs for audio scene modeling, the research by [chen2018deep] demonstrated better performance for the model trained with scalograms compared to one using spectrograms.
@@
As illustrated above, while numerous studies have explored audio recognition using spectrograms or scalograms in conjunction with deep learning, few have directly compared the performance of these two feature extraction methods. Existing benchmarks often offer indirect comparisons, as they involve different prediction models rather than a comprehensive evaluation of the feature extractors themselves, as seen in the study by [chen2018deep]. Or the study by [huzaifah2017comparison] compared different transforms but focused solely on predictive performance, lacking a comparative analysis of how the attributes of the transforms influence predictive performance and computational expense. The research by [tzanetakis2001audio] compared various time-frequency transforms but used discrete coefficients directly instead of scalograms or spectrograms, limiting the models’ performance. Given these limitations, this paper aims to develop an approach for a rigorous comparison of spectrograms and scalograms as audio feature extractors. To this end, the experimental design for both methods is kept identical, including the original audio data used, the format of input data, the configuration of the CNNs models, and the training and evaluation methods. The sole distinction in this study is the type of feature extractor: spectrogram versus scalogram.
@@ II-A Short-time Fourier transform
The STFT [muller2015fundamentals] is a method used to analyze signals by dividing them into segments based on time frames and performing Fourier transforms (FT) on each segment, denoted as (1). This approach allows the signal to be examined in both the time shift $\tau$ and frequency shift $\omega$ simultaneously. By applying a windowing function, denoted as $\gamma(t)$, to isolate specific portions of the signal, STFT computes the spectrum over short durations, enabling it to capture not only the frequencies present in the signal but also how these frequencies change over time.
$X_{STFT}(\tau,\omega)=\int_{-\infty}^{\infty}\,x(t)\gamma^{*}(t-\tau)e^{-j\omega t}\,dt$ (1)
@@
In practical applications, signals are processed in discrete time formats, and the STFT is computed using a window of length N, referred to as the frame size, denoted as (2). Since shifting sample-wise results in redundant data due to the similarity of neighboring spectra, a larger step size, known as the hop size H, is used, typically set to half the window length H=N/2. This choice strikes a balance between achieving high time resolution and managing data volume effectively [muller2015fundamentals]. Each sample point in the data sequence, where STFT is calculated, corresponds to a frame index m defined as m=M/H, with M representing the total data length. Within each frame, computing the FT of a discrete signal of N points in time results in a frequency vector of N points. Due to the symmetric nature of the Fourier spectrum, only the first half of the spectrum (k=0 to N/2) is typically considered to avoid redundancy. Consequently, the output of STFT is structured as a matrix with dimensions (M/H, N/2+1) where each element X(m,k) represents the complex Fourier coefficient of the $m^{th}$ time frame at frequency index k.
To enhance the visualization of STFT results, spectrograms are commonly employed [mertins1999signal]. Represented as a type of heatmap image, they provide a two-dimensional representation of signal energy across time and frequency dimensions. Time is arranged along the x-axis, and frequency along the y-axis. The intensity of color in each pixel of the spectrogram corresponds to the degree of signal energy. For instance, the colors range from dark blue, indicating lower energy, to dark red, indicating higher energy, as illustrated in Fig. 1(a).
@@ II-B Wavelet transform
$X_{WT}(\textit{b},\textit{a})=\frac{1}{\sqrt{\textit{a}}}\int_{-\infty}^{\infty}\,x(t)\psi^{*}(\frac{t-\textit{b}}{\textit{a}})\,dt$ (3)
@@ II-C Time and frequency resolution of transforms
To achieve a comprehensive analysis of a signal, both time resolution and frequency resolution should be maximized. This implies capturing signal variations with minimal shifts in both time and frequency. However, the uncertainty principle [mertins1999signal] imposes a constraint on the tradeoff between time and frequency resolution, as described in (5). Utilizing a narrow time window (small $\Delta_{t}$) allows for high time resolution in STFT or WT, but results in poor frequency resolution (large $\Delta_{\omega}$). Conversely, employing a large time window yields poor time resolution but improved frequency resolution.
@@ III-A Workflow
The entire workflow of the experiment in this study is presented in Fig. 3. Initially, the audio data is normalized to constrain the amplitude values within the range (0, 1) before being transformed using the STFT and WT, respectively. Following the transformations, the matrices produced as outputs from the STFT and WT are used to create spectrograms and scalograms. These are then plotted as images and split into training and validation sets for a CNNs model. The performance of the trained CNNs is then recorded and compared between the two types of data, using metric the Area Under the Curve of the Receiver Operating Characteristic (AUC-ROC) [provost1998case].
@@ III-B MIMII audio dataset
The acoustic recognition task in this study is conducted using the MIMII audio dataset for machinery fault detection [purohit2019mimii]. This dataset was created to provide real-life sound data from factories for various machine types, including fans, pumps, sliders, and valves. These machines produce both stationary and non-stationary sounds, presenting varying levels of difficulty in distinguishing between normal and abnormal conditions. Background noise from various factory environments is recorded and mixed with the target machine sounds to simulate real-world conditions at three levels of signal-to-noise ratio (SNR): -6 dB, 0 dB, and 6 dB. Each audio recording has a length of 10 seconds, sampled at 16 kHz, resulting in each discrete audio 160,000 samples.
@@ III-D Implementation of Short-Time Fourier Transform
The STFT is conducted in a time-discrete manner using Librosa [mcfee2015librosa]. Initially, the frame size N and hop size H are treated as hyperparameters aimed at optimizing the performance of the trained model during the initial stages of training. Subsequently, a configuration of STFT with N = 1024 and H = 512 demonstrating good predictive performance is selected, resulting a matrix with dimensions (513, 313). This configuration is then applied uniformly across the entire audio dataset to generate the STFT representations.
@@ III-E Implementation of Continuous Wavelet Transform
The CWT is also executed in a time-discrete manner using Pywavelets [lee2019pywavelets]. The translation is performed sample-wise, and the scale parameter encompasses a range of consecutive natural numbers. Given that the scale determines the frequency resolution, various scales are evaluated during the training of the CNNs model. Then the scale from 2 to 129, demonstrating optimal predictive performance, is selected, producing output matrices with dimensions (16000, 128). Subsequently, this selected scale is employed to generate scalograms for the entire audio dataset.
@@ III-F Implementation of Convolutional Neural Networks
The primary focus of this study is to compare the performance between spectrograms and scalograms. Therefore, the CNNs model is designed in an exploratory manner, ensuring it achieves satisfactory performance. The Tensorflow [tensorflow2015-whitepaper] is used for CNNs implementation. Several configurations of CNNs are evaluated, and an effective configuration is chosen based on its performance, as outlined in Table I.
@@ IV-A Benchmark study
The benchmark study [gantert2021supervised], representing the state-of-the-art for supervised methods on the same dataset, employed Mel-frequency cepstral coefficients (MFCCs) as spectral features. Unlike the approach of transforming data into compressed images, this model directly utilized tuples generated by MFCCs. Prior to being used for training, the tuples were normalized to have zero mean and unit standard deviation. The prediction model used in this study was a Multilayer Perceptron (MLP). Additionally, the tuples of abnormal sounds were oversampled to achieve a balanced dataset. Detailed comparisons are presented in Table II.
@@ IV-B Performance evaluation
For the comparison between spectrogram and scalogram within each machine type, the results are presented in Table V and visualized in Figure 6. Overall, spectrograms consistently outperform scalograms, with the exception of one case involving valves where CNNs predicts audio based on scalograms more effectively than those based on spectrograms. This exception arises due to the non-stationary nature of valve audio signals [purohit2019mimii], which are impulsive and sparse in time. The WT used in scalogram generation leverages its multiresolution feature to enhance feature extraction for such non-stationary signals [mertins1999signal]. Conversely, when evaluating performance on fan audio signals, spectrograms achieve significantly higher performance compared to scalograms. This difference can be attributed to the stationary nature of fan sounds [purohit2019mimii], where the constant time and frequency resolution of spectrograms are more adept at feature extraction than scalograms.